# ELIP: Enhanced Visual-Language Foundation Models for Image Retrieval

### v2

@@ Abstract
Figure 1: The ELIP architecture. Left: We propose a novel architecture that can be applied to pre-trained and frozen vision-language foundation models, such as CLIP, SigLIP, SigLIP-2 and BLIP-2, to enhance their text-to-image retrieval performance. The key idea is to use the text query to define a set of visual prompt vectors that are incorporated into the image encoder to make it aware of the query when generating the embedding. An MLP maps from the text space to the visual space of the input to the ViT encoder. The architecture is lightweight, and our data curation strategies enable efficient and effective training with limited resources. Right: In this retrieval example from the COCO benchmark, the top-$k$ ($k$=100) images are re-ranked by our ELIP model for the text query: ‘People on bicycles ride down a busy street’. The ground truth image matching the query is not in the top-5 ranked images in the initial CLIP ranking, but is ranked top-1 (highlighted in the dashed box) by the re-ranking.
@@ 2 Related Work
Text-to-Image Retrieval is a fundamental and much researched task in cross-modal learning [46, 13, 40, 26, 102, 68, 47, 16, 86, 97, 17, 12, 14, 15, 20, 21, 23, 30, 36, 39, 43, 44, 49, 52, 53, 75, 78, 84, 85, 87, 88, 90, 91, 95, 98, 101, 103, 104, 105]. Large vision language models such as CLIP [68, 37], ALIGN [41], BLIP-2 [48], SigLIP [99] and SigLIP-2 [81] that have powerful zero-shot capabilities have now become the de facto method for open-set text-based image retrieval. The most recent work [74] gives a slight improvement over BLIP-2 by incorporating the output of an object detector or annotations of detection bounding boxes. This succeeds in overcoming the failure cases where small but semantically important objects in an image are not properly understood by the model. We compare to this model and show superior performance.
@@ 4.1 Text-Guided MLP Mapping Network
where $T$ denotes the query text, which is first encoded with a pre-trained, frozen text encoder ($\Phi_{\text{t}}(\cdot)$) into $m+1$ embeddings. The [CLS] token is further fed into a trainable mapping network to generate the prompt vectors, which are concatenated with the $n+1$ image embeddings ($[x_{p}^{1},\dots,x_{p}^{n},x_{\texttt{CLS}}]$), and passed into the pre-trained, frozen visual encoder ($\Phi_{\text{v}}(\cdot)$). The MLP Mapping Network consists of 3 layers of linear layers with a GELU between every two linear layers. The ELIP architecture is shown in Figure 2 and Figure 3.
@@ 4.3 Training and Testing ELIP-B
Text-Guided Image-Text Matching Loss. At training time, we feed the text query ($T$) and the re-computed image features with the query text as prompts, i.e.,$\{v_{+},v_{-}\}$ ($v_{+}$ denotes the positive image and $v_{-}$ denotes the negative image), into the Q-Former, and then to an Image-Text Matching (ITM) Head to predict a score indicating whether the text and image match or not. The output of the ITM head is trained with binary cross entropy loss.
@@ 5.2 Selection and Curation of Large-Scale Dataset
Setting Architecture Training Dataset Hard Sample Mining Multiple Prompts COCO Avg. Flickr Avg.
@@ 6.2 Out-of-Distribution Benchmarks
Occluded COCO is curated with annotations from [45], with the method as described in [100], where the occlusion relationship is utilised to collect images containing occluded objects. This dataset aims to evaluate the model’s performance on retrieving images with occluded target objects against images that do not contain the target object.
ImageNet-R is generated using annotations from [34] and aims to examine the model’s performance for retrieval across various domains, for example, art, cartoons, deviantart, graffiti, embroidery, graphics, origami, paintings, patterns, plastic objects, plush objects, sculptures, sketches, tattoos, toys, and video games.
@@ 7.1 Results on COCO and Flickr Benchmarks
Comparison with State-of-the-Art. As shown in Table 2, we compare our models (ELIP-C, ELIP-S, ELIP-S-2, and ELIP-B) with prior state-of-the-art methods. When trained on DataCompDR12M, our method demonstrates zero-shot performance improvements for CLIP, SigLIP, SigLIP-2, and BLIP-2 on the COCO and Flickr benchmarks. Notably, ELIP-B outperforms the most recent work [74], establishing a new state-of-the-art for text-to-image retrieval on the BLIP-2 backbone. Furthermore, our ELIP-S, when applied to SigLIP and SigLIP-2, achieves performance comparable to BLIP-2.
@@
Model Occluded COCO ImageNet-R Average
@@ A.3 Implementation Details for Fine-Tuning on COCO and ImageNet
As mentioned in Section 7.2, we have further fine-tuned our models (pre-trained on DataCompDR) on COCO and ImageNet, respectively. For both datasets, we first group the hard samples in batches using the method described in Section 5.1. The captions for the images are category names, and we ensure that in a batch one category can only appear once, so that there will not be noise in training. We fine-tune ELIP-C, ELIP-S, and ELIP-B with an initial learning rate of $51e-$ 5 for 10k iterations. Although there is still a domain gap between COCO and Occluded COCO, ImageNet, and ImageNet-R, we find that even fine-tuning on COCO and ImageNet can already bring a boost to the performance on Occluded COCO and ImageNet-R as the caption domain is closer. It is not feasible to fine-tune on Occluded COCO and ImageNet-R because Occluded COCO has very few data samples and ImageNet-R is designed for evaluation.
@@ A.4 Implementation Details for Setting up the Occluded COCO Benchmark
Appendix B Additional Qualitative Results
@@
B.1 Qualitative Comparison of CLIP and ELIP-C
@@
B.2 Qualitative Comparison of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
@@
B.3 Qualitative Comparison of BLIP and ELIP-B
@@
B.4 Attention Map of CLIP and ELIP-C
@@
B.5 Attention Map of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
@@
B.6 Attention Map of BLIP-2 and ELIP-B
@@
Appendix C Ablation Study for Different Number of Generated Visual Prompt Tokens
@@
Appendix D Further Ablation Study for ELIP-B
@@
Appendix E Efficiency of ELIP Pre-Training
      Model       GPU Hours       #GPU       BS       FLOPS       $\Delta$FLOPS

### v3

@@ Abstract
Figure 1: The ELIP architecture. Left: We propose a novel architecture that can be applied to pre-trained and frozen vision-language foundation models, such as CLIP, SigLIP, SigLIP-2 and BLIP-2, to enhance their text-to-image retrieval performance. The key idea is to use the text query to define a set of visual prompt vectors that are incorporated into the image encoder to make it aware of the query when generating the embedding. An MLP maps from the text space to the visual space of the input to the ViT encoder. The architecture is lightweight, and our data curation strategies enable efficient and effective training with limited resources. Right: In this retrieval example from the COCO benchmark, the top-$k$ (kk=100) images are re-ranked by our ELIP model for the text query: ‘People on bicycles ride down a busy street’. The ground truth image matching the query is not in the top-5 ranked images in the initial CLIP ranking, but is ranked top-1 (highlighted in the dashed box) by the re-ranking.
@@ 2 Related Work
Text-to-Image Retrieval is a fundamental and much researched task in cross-modal learning [47, 15, 104, 69, 48, 18, 88, 99, 19, 16, 17, 22, 23, 25, 31, 37, 40, 43, 44, 50, 53, 54, 76, 79, 86, 87, 89, 90, 92, 93, 97, 100, 103, 105, 106, 107, 84, 45]. Large vision language models such as CLIP [69, 38], ALIGN [41], BLIP-2 [49], SigLIP [101] and SigLIP-2 [82] that have powerful zero-shot capabilities have now become the de facto method for open-set text-based image retrieval. The most recent work [75] gives a slight improvement over BLIP-2 by incorporating the output of an object detector or annotations of detection bounding boxes. This succeeds in overcoming the failure cases where small but semantically important objects in an image are not properly understood by the model. We compare to this model and show superior performance.
@@ 4.1 Text-Guided MLP Mapping Network
where $T$ denotes the query text, which is first encoded with a pre-trained, frozen text encoder ($\Phi_{\text{t}}(\cdot)$) into $m+1$ embeddings. The [CLS] token is further fed into a trainable mapping network to generate the prompt vectors, which are concatenated with the $n+1$ image embeddings ($x_{p}^{1},\dots,x_{p}^{n},x_{\texttt{CLS}}]$), and passed into the pre-trained, frozen visual encoder ($\Phi_{\text{v}}(\cdot)$). The MLP Mapping Network consists of 3 layers of linear layers with a GELU between every two linear layers. We expand the output dimension to be $n$ times when we generate $n$ tokens and then divide the generated vector into $n$ tokens. The ELIP architecture is shown in Figure 2 and Figure 3.
@@ 4.3 Training and Testing ELIP-B
[…]
@@ 5.2 Selection and Curation of Large-Scale Dataset
Setting
Architecture
Training
Dataset
Hard Sample
Mining
Multiple
Prompts
COCO Avg. Flickr Avg.
@@ 6.2 Out-of-Distribution Benchmarks
Occluded COCO is curated with annotations from [46], with the method as described in [102], where the occlusion relationship is utilised to collect images containing occluded objects. This dataset aims to evaluate the model’s performance on retrieving images with occluded target objects against images that do not contain the target object. It has 76 text queries and 5,000 images.
ImageNet-R is generated using annotations from [35] and aims to examine the model’s performance for retrieval across various domains, for example, art, cartoons, deviantart, graffiti, embroidery, graphics, origami, paintings, patterns, plastic objects, plush objects, sculptures, sketches, tattoos, toys, and video games. It has 200 text queries and 30,000 images.
@@ 7.1 Results on COCO and Flickr Benchmarks
Comparison with State-of-the-Art. As shown in Table 2, we compare our models (ELIP-C, ELIP-S, ELIP-S-2, and ELIP-B) with prior state-of-the-art methods. When trained on DataCompDR12M, our method demonstrates zero-shot performance improvements for CLIP, SigLIP, SigLIP-2, and BLIP-2 on the COCO and Flickr benchmarks. Notably, ELIP-B outperforms the most recent work [75], establishing a new state-of-the-art for text-to-image retrieval on the BLIP-2 backbone. Furthermore, our ELIP-S, when applied to SigLIP and SigLIP-2, achieves performance comparable to BLIP-2. We have also compared ELIP with several baseline methods for re-ranking in the appendix.
@@
Model
Occluded
COCO
ImageNet-R Average
@@ A.3 Implementation Details for Fine-Tuning on COCO and ImageNet
As mentioned in Section 7.2, we have further fine-tuned our models (pre-trained on DataCompDR) on COCO and ImageNet, respectively. For both datasets, we first group the hard samples in batches using the method described in Section 5.1. The captions for the images are category names, and we ensure that in a batch one category can only appear once, so that there will not be noise in training. We fine-tune ELIP-C, ELIP-S, and ELIP-B with an initial learning rate of 1e51e-5 for 10k iterations. Although there is still a domain gap between COCO and Occluded COCO, ImageNet, and ImageNet-R, we find that even fine-tuning on COCO and ImageNet can already bring a boost to the performance on Occluded COCO and ImageNet-R as the caption domain is closer. It is not feasible to fine-tune on Occluded COCO and ImageNet-R because Occluded COCO has very few data samples and ImageNet-R is designed for evaluation.
@@ A.4 Implementation Details for Setting up the Occluded COCO Benchmark
Appendix B Model Variations and Comparison with Other Re-Ranking Methods
Model ID Recall@1 Recall@5 Recall@10
CLIP - 67.56 88.34 93.00
Late Fusion Baseline 1 69.58 89.58 93.66
Dense Text Token 2 71.66 90.44 94.54
Cross-Attention Pooling 3 67.78 88.42 93.52
Text Query Expansion 4 66.28 87.60 92.88
Image Crop Features 5 67.52 88.36 92.98
QE + LF 6 67.18 88.44 93.00
ELIP-C (ours) - 72.30 90.62 94.68
Table 5: Model variations for ELIP-C (ID 1-3), and comparison with other re-ranking methods (ID 4-6). The performance is measured on Flickr retrieval. See the text for a detailed description of each method.
In this section, we investigate variations of the ELIP model and also compare with several other re-ranking methods as mentioned in Section VI of the main paper. We report results using CLIP on Flickr.
Variations on the ELIP model
  • 1. Late Fusion Baseline: For both training and inference, the MLP Mapping Network is used to map the text embedding to generate multiple prompt vectors that are inserted into the last layer of the ViT image encoder rather than the first layer.
  • 2. Dense Text Token: Rather than use the $t_{CLS}$ embedding to generate all prompt vectors, for both training and inference, we use the average pooling of dense text token representations to generate the prompt vectors.
  • 3. Cross-Attention Pooling: Given the image patch embeddings at the final ViT layer, for both training and inference, we map the $t_{CLS}$ vector to the same dimension as the image patch tokens via a MLP network, and use it to cross-attend to the image patch embeddings to get the image feature, instead of generating prompt vectors and inserting into ViT image encoder. The idea of cross-attention pooling is similar to [7].
Comparisons with other methods
  • 4. Text Query Expansion: For each text query, we ask GPT-4 [3] to rephrase it into 5 variants. At the re-ranking stage, we compute the similarity scores of the top-$k$ images with the 5 variants of text queries, and then take an average of the 5 scores together with the initial score to generate the final score.
  • 5. Image Crop Features: For each of the top-$k$ images, we randomly crop and resize it into 5 variants and feed them into the CLIP image encoder to get the image local features. Then we compute the similarity scores between the image local features and the text features, and take an average of the 5 scores together with the initial score to generate the final score.
  • 6. Query Expansion + Local Features: We use the top-3 retrieved images as expanded query [20] for re-ranking, i.e., the top-3 retrieved image features are used as query features in the re-ranking. We then compare the Chamfer Similarity [8, 45] between patch features of the expanded query images and other top-$k$ images. The patch features are the patch tokens of the ViT image encoder. The average of the Chamfer similarities and the initial score is used as the final score for re-ranking. We have also tried to directly match image patch tokens with text dense tokens, but the performance is not good due to the fact that text tokens and image patch tokens are not properly aligned.
It can be observed that the ELIP-C model achieves the best performance over the variations and in comparison to other methods.
Appendix C Additional Qualitative Results
@@
C.1 Qualitative Comparison of CLIP and ELIP-C
@@
C.2 Qualitative Comparison of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
@@
C.3 Qualitative Comparison of BLIP and ELIP-B
@@
C.4 Attention Map of CLIP and ELIP-C
@@
C.5 Attention Map of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
@@
C.6 Attention Map of BLIP-2 and ELIP-B
@@
Appendix D Ablation Study for Different Number of Generated Visual Prompt Tokens
@@
Appendix E Further Ablation Study for ELIP-B
@@
Appendix F Efficiency of ELIP Pre-Training
[…]